# Relative energy constraint in Boltzmann’s surface entropy eliminates thermodynamic inconsistencies and allows negative absolute temperatures

### v1

@@ Abstract
Over the past century, an intense debate in statistical mechanics has been about the correctness of Boltzmann’s surface entropy versus Gibbs’ volume entropy, for isolated systems. Both entropies make significantly different predictions for systems with few degrees of freedom. Even in the thermodynamic limit, they can disagree—while Boltzmann entropy allows negative absolute temperatures to exist, Gibbs entropy precludes such a possibility. Here, we show that modifying Boltzmann’s entropy via a relative energy tolerance eliminates thermodynamic inconsistencies in several model systems with unbounded energy spectra by ensuring positive, finite temperatures. Concomitantly, the proposed entropy allows for negative temperatures in systems with bounded spectra and closely matches canonical ensemble predictions. This work conclusively remedies the prevalent deficiencies of the Gibbs and Boltzmann entropy formulations and paves the way for the use of the modified Boltzmann entropy in the microcanonical ensemble, allowing negative temperatures to exist.
@@
The microcanonical ensemble in statistical mechanics, originally proposed by Gibbs over a century ago [1], continues to receive a significant extent of attention [2, 3, 4, 5, 6]. In the physics literature, there exist two competing entropy definitions in the microcanonical ensemble: Gibbs’ volume entropy (also referred to as the Hertz entropy) [1, 7, 8, 9, 10], based on the phase space volume and Boltzmann’s surface entropy [11, 12] (also referred to as the Boltzmann-Planck entropy), based on the phase space density. Even though Boltzmann’s surface entropy dominated textbook discourses [13, 14, 15, 16, 17], many studies favored the use of the Gibbs entropy [18, 8, 19, 9, 10, 5]. About a decade ago, Dunkel and Hilbert indicated that Boltzmann’s surface entropy leads to violations of equipartition and unphysical results, i.e., negative/infinite temperatures, for some model systems with unbounded spectra [5]. Subsequently, several articles appeared that pointed out that Gibbs entropy suffers from serious objections such as the consideration of all states with energy lower than energy $E$, rather than only the states at the desired energy, and the inability to allow for negative temperatures in systems with non-monotonically increasing density of states [20, 12, 21, 22, 23, 24, 6].
@@
Although the Boltzmann and Gibbs entropies usually agree for very large systems, they differ significantly and lead to widely varying temperatures for isolated systems with few degrees of freedom [25]. (In some cases, even in the thermodynamic limit, e.g., the case of a two-state system (see below), they can disagree.) Note that the principles of statistical mechanics can be applied to systems with few degrees of freedom, as the statistical averaging is carried out over the classical phase space or the available quantum states, rather than over the degrees of freedom [26, 27, 28, 6]. Thus, even quantities like the temperature of an isolated quantum state [29, 30, 31] or the entropy of a few-body system are well defined [32, 33].
@@
In light of this fact, that both entropy definitions have shortcomings for small systems [22] is particularly unnerving. Indeed, the lack of a universal microcanonical entropy formulation not only adversely impacts the fundamental physical understanding of the thermodynamics of small systems [34], but also could impede technological progress in the future, wherein systems of few atoms, molecules, or electrons are increasingly becoming physically realizable [35]. In this work, we show that the consideration of a relative energy tolerance in the microcanonical ensemble eliminates all thermodynamic inconsistencies put forth in previous work pertaining to systems with unbounded spectra, while at the same time allowing negative temperatures in systems with bounded spectra.
@@
Formulation of the relative-energy constrained Boltzmann entropy. Considering the microcanonical ($NVE$) ensemble with fixed number of particles $N$, volume $V$, and energy $E$, the Gibbs entropy ($S_{G}$) is defined as $S_{G}=k_{B}\ln\Omega$, where $\Omega$ denotes the volume of the classical phase space or the number of quantum states in the region $E^{\prime}<E$ (Figure 1a). For a classical system, this can be written as
@@
where $\Theta$ is the Heaviside step function, $\Theta(x)=\begin{cases}1;\;\;\;x>0\\ 0;\;\;\;x\leq 0\end{cases}$, $h$ is Planck’s constant, $d$ is the dimensionality of the system, $\mathrm{\textbf{r}}^{N}$ denotes the positions of the particles, and $\mathrm{\textbf{p}}^{N}$ denotes their momenta. Alternatively, for a quantum-mechanical system, we can write
@@
where $\mathrm{Tr}$ denotes the trace, i.e., the sum over quantum states in a Hilbert space. In contrast, the Boltzmann entropy ($S_{B}$) is conventionally defined as $S_{B}=k_{B}\ln\omega$, where
@@
and $\epsilon$ is an absolute energy tolerance, as visualized in Figure 1a. Physically, the tolerance appears because no system is perfectly isolated, and some amount of perturbation or broadening in the energy occurs due to the system’s interactions with its surroundings (Figure 1b).
From Eq. (1) and (2), one can infer that
@@
Even for a discrete spectrum, although direct enumeration is often used as a simplification [36], Eq. (3) is preferred [14, 37], as it allows the correct definition of entropy for nondegenerate quantum systems [38, 39]. Typically, $\epsilon$ has been interpreted to be a fixed constant in Eq. (3), and this has led to inconsistent results for simple systems with unbounded energy levels, e.g., negative temperatures for an ideal gas with one degree of freedom or infinite temperature for a quantum harmonic oscillator and an ideal gas with two degrees of freedom. We suggest here that instead of using an absolute energy tolerance, the energy tolerance in the microcanonical ensemble should be defined relative to the system’s energy (Figure 1c). Indeed, a system with 1 J of energy and with 1 eV = 1.602 $\times$ 10-19 J of energy should not use the same tolerance in the microcanonical ensemble. This is because the higher the energy of a system, the greater its interaction with its surroundings, causing more broadening or perturbation in its energy. Thus, we propose that the constraint inside the delta function should be expressed in terms of the relative error of $H_{r}=\frac{H}{E}$ from 1. In terms of the phase space integrals or Hilbert space summations, doing so amounts to applying a constraint on $H$ without using up a degree of freedom. Thus, classically, one should write
@@
where $\epsilon_{r}$ is a relative error tolerance. In Eqs. (4) and (5), we have used the subscript $r$ to denote the use of the relative deviation from the desired energy. It can be shown that $\omega_{r}=\epsilon_{r}E\frac{\partial\Omega}{\partial E}$, as indicated in Figure 1c. Note that there is some precedent in the literature for adopting an energy-dependent tolerance [39, 38].
We now apply the new approach for calculating the microcanonical relative-energy-constrained Boltzmann entropy, $S_{B_{r}}=k_{B}\ln\omega_{r}$, and the corresponding temperature
@@
Table 1: Comparison of the predictions of the Gibbs, conventional Boltzmann, and proposed modified Boltzmann entropies in the microcanonical ensemble with the corresponding predictions in the canonical ensemble. Only the relative-energy-constrained Boltzmann entropy leads to predictions consistent with the canonical ensemble in every case considered.
System Gibbs Conventional Boltzmann Modified Boltzmann Canonical
3D ideal gas $E=\frac{3N}{2}k_{B}T$ $E=\left(\frac{3N}{2}-1\right)k_{B}T$ $E=\frac{3N}{2}k_{B}T$ $E=\frac{3N}{2}k_{B}T$
Harmonic oscillator (high $T$) $E=k_{B}T-\frac{h\nu}{2}$ $T=\infty$ $E=k_{B}T$ $E=k_{B}T$
Harmonic oscillator (any $T$) [^1] Note that $E_{0}=\frac{1}{2}h\nu$ and $\theta=\frac{h\nu}{k_{B}}$ for the harmonic oscillator. $\frac{E-E_{0}}{h\nu}=\frac{3-\exp\left(\frac{\theta}{T}\right)}{2\exp\left(\frac{\theta}{T}\right)-2}$ $T=\infty$ $\frac{E-E_{0}}{h\nu}=\frac{1}{\exp\left(\frac{\theta}{T}\right)-1}$ $\frac{E-E_{0}}{h\nu}=\frac{1}{\exp\left(\frac{\theta}{T}\right)-1}$
Particle in 1D box $E=\frac{k_{B}T}{2}$ $E=-\frac{k_{B}T}{2}$ $E=\frac{k_{B}T}{2}$ $E=\frac{k_{B}T}{2}$
Negative temperatures? No Yes Yes Yes
Application of the modified entropy to example systems: $d$-dimensional monoatomic ideal gas. Consider a $d$-dimensional monoatomic ideal gas consisting of $N$ atoms of mass $m$ each, occupying a total volume $V$. This system has unbounded energy levels and thus should not have negative temperature, as correctly argued by Dunkel and Hilbert[5]. The volume of the phase space for this system is [19]
$\Omega=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}+1\right)}E^{\frac{dN}{2}}.$
@@
$\omega=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}\right)}E^{\frac{dN}{2}-1}\epsilon.$
@@
The Gibbs temperature ($T_{G}$) can be calculated using the equation $\frac{1}{T_{G}}=\frac{\partial S_{G}}{\partial E}$ to obtain $E=\frac{dN}{2}k_{B}T_{G}$. However, using the analogous equation for Boltzmann temperature, assuming a fixed energy tolerance $\epsilon$, leads to $E=\left(\frac{dN}{2}-1\right)k_{B}T_{B}$. Note that if $dN=1$, i.e., the system has one degree of freedom, one obtains negative temperatures whereas if $dN=2$, i.e., the system has two degrees of freedom, one gets infinite temperature [5]. This is a serious shortcoming of the Boltzmann entropy, as pointed out by Dunkel and Hilbert [5], as one is led to predict negative/infinite temperature for a system whose energy is unbounded. Considering a fractional energy tolerance, one however obtains $\omega_{r}=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}\right)}E^{\frac{dN}{2}}\epsilon_{r}$, which combined with Eq. (6) leads to
$E=\frac{dN}{2}k_{B}T_{B_{r}}.$
The above equation not only disallows negative/infinite temperatures for an ideal gas but fixes the disagreement between the energy of a three-dimensional ideal gas as traditionally calculated in the canonical $\left(E=\frac{3N}{2}k_{B}T\right)$ and microcanonical $\left(E=\left(\frac{3N}{2}-1\right)k_{B}T\right)$ ensembles (see Table 1)!

### v2

# Resolving the Debate Between Boltzmann and Gibbs Entropy: Relative Energy Window Eliminates Thermodynamic Inconsistencies and Allows Negative Absolute Temperatures

@@ Abstract
Small systems consisting of a few particles are increasingly technologically relevant. In such systems, an intense debate in microcanonical statistical mechanics has been about the correctness of Boltzmann’s surface entropy versus Gibbs’ volume entropy. Both entropies have shortcomings—while Boltzmann entropy predicts unphysical negative/infinite absolute temperatures for small systems with unbounded energy, Gibbs entropy entirely disallows negative absolute temperatures, in disagreement with experiments. We consider a relative energy window, motivated by the Heisenberg energy-time uncertainty principle and eigenstate thermalization in quantum mechanics. The proposed entropy ensures positive, finite temperatures for systems without a maximum limit on their energy and allows negative absolute temperatures in bounded energy systems, e.g., with population inversion. It also closely matches canonical ensemble predictions for prototypical systems, thus correctly describing the zero-point energy of an isolated quantum harmonic oscillator. Overall, we enable accurate thermodynamic models for isolated systems with few degrees of freedom.
IISC]Department of Chemical Engineering, Indian Institute of Science, Bengaluru, Karnataka 560012, India
{tocentry}[Uncaptioned image]
@@
The microcanonical ensemble in statistical mechanics, originally proposed by Gibbs over a century ago 1, continues to receive a significant extent of attention 2, 3, 4, 5, 6, 7. In the physics literature, there exist two competing entropy definitions in the microcanonical ensemble: Gibbs’ volume entropy (also referred to as the Hertz entropy) 1, 8, 9, 10, 11, based on the phase space volume and Boltzmann’s surface entropy 12, 13 (also referred to as the Boltzmann-Planck entropy), based on the phase space density. Even though Boltzmann’s surface entropy dominated textbook discourses 14, 15, 16, 17, 18, many studies favored the use of the Gibbs entropy 19, 9, 20, 10, 11, 5. Note that the terms Gibbs entropy for the definition based on the phase space volume and Boltzmann entropy for that based on the phase space density are based on the prevalent names in the literature, although Gibbs’ seminal work did acknowledge the possibility of both entropy definitions1.
About a decade ago, Dunkel and Hilbert indicated that Boltzmann’s surface entropy leads to violations of equipartition and unphysical results, i.e., negative/infinite temperatures, for some model systems with unbounded spectra 5, i.e., those without a maximum limit on the allowable energies. Subsequently, several articles appeared that pointed out that Gibbs entropy suffers from serious objections, such as the consideration of all states with energy lower than energy $E$, rather than only the states at the desired energy and the inability to allow for negative absolute temperatures in systems with non-monotonically increasing density of states 21, 13, 22, 23, 24, 25, 6. Physically, a negative absolute temperature implies that the entropy of a system reduces as the energy increases. This is indeed possible in systems with bounded energy spectra as higher energies would involve more particles occupying the highest-energy level, thus reducing the entropy with an increase in energy.
@@
Although the Boltzmann and Gibbs entropies usually agree for very large systems, they differ significantly and lead to widely varying temperatures for isolated systems with few degrees of freedom 26. (In some cases, even in the thermodynamic limit, e.g., the case of a two-level system (see below), they can disagree.) In this regard, the principles of statistical mechanics can be applied to systems with few degrees of freedom, as the statistical averaging is carried out over the classical phase space or the available quantum states, rather than over the degrees of freedom 27, 28, 29, 6. Thus, even quantities like the temperature of an isolated quantum state 30, 31, 32 or the entropy of a few-body system are well defined 33, 34.
@@
In light of this fact, that both entropy definitions have shortcomings for small systems 23 is particularly unnerving. Indeed, the lack of a universal microcanonical entropy formulation not only adversely impacts the fundamental physical understanding of the thermodynamics of small systems 35, but also could impede technological progress in the future, wherein systems of few atoms, molecules, or electrons are increasingly becoming physically realizable 36. Not only that, the applicability of statistical mechanics to small systems is particularly important in molecular biology37, nuclear physics,38, 39 and astronomical systems40. In fact, Puglisi et al. have discussed that the conceptual framework of statistical mechanics can successfully predict the behavior of small systems.41
In this Letter, we show that the consideration of a relative energy tolerance (also called the energy window) in the microcanonical ensemble eliminates all thermodynamic inconsistencies put forth in previous work pertaining to systems with unbounded spectra while at the same time allowing negative absolute temperatures in systems with bounded spectra. In addition, in the Supporting Information, we show that a consistency condition mentioned in ref.  5 is not valid even for the Gibbs entropy, thus invalidating the former’s use to support the latter. Now, by the term unbounded, we refer to the fact that the maximum allowable system energy is infinite, i.e., it has no upper bound. Conversely, the term bounded means that the system energy has an upper limit, as seen in the case of a two-level spin system later in the paper. Note that the term “bounded” does not have any reference to the energy constraint employed in Boltzmann-type entropies.
It is also important to clarify that our focus is not on the entropy formulae ascribed to Boltzmann and Gibbs, which are both equally valid. Rather, we seek to examine the controversy surrounding the correct choice of the microstates that are counted to determine the entropy in Boltzmann’s entropy formula, which is applicable to a microcanonical system as
$S=k_{B}\ln w$ (1)
where $S$ denotes the entropy of the system, $k_{B}$ is the Boltzmann constant, and $w$ is the number of microstates, each having equal probability $1/w$. In the above formula, Gibbs’ entropy uses the phase space volume ($\Omega$) for $w$, and Boltzmann’s entropy uses the phase space density ($\omega$) for it. Note that Gibbs’ entropy is unrelated to the Gibbs entropy formula, which allows one to calculate entropy in any statistical-mechanical ensemble as:
$S=-k_{B}\sum_{i}p_{i}\ln p_{i}$ (2)
where $p_{i}$ denotes the probability of occurrence of the microstate $i$, the summation running over all possible microstates corresponding to the considered macrostate.
@@
Now, considering the microcanonical ($NVE$) ensemble with fixed number of particles $N$, volume $V$, and energy $E$, the Gibbs entropy ($S_{G}$) is defined as $S_{G}=k_{B}\ln\Omega$, where $\Omega$ denotes the volume of the classical phase space or the number of quantum states in the region $E^{\prime}<E$ (Figure 1a), and $E^{\prime}$ denotes the possible values the system’s energy can take. For a classical system, this can be written as
@@
where $\Theta$ is the Heaviside step function, $\Theta(x)=\begin{cases}1;\;\;\;x>0\\ 0;\;\;\;x\leq 0\end{cases}$, $h$ is Planck’s constant, $d$ is the dimensionality of the system, $H$ is the Hamiltonian of the system, $\mathrm{\textbf{r}}^{N}$ denotes the positions of the particles, and $\mathrm{\textbf{p}}^{N}$ denotes their momenta. Alternatively, for a quantum-mechanical system, we can write
@@
where $\mathrm{Tr}$ denotes the trace, i.e., the sum over quantum states in a Hilbert space. The ensemble corresponding to this entropy definition has sometimes been referred to as the uniform ensemble, where the probability of every state with $H\leq E$ is equal,42, 16 rather than only those with $H=E$ as in the microcanonical ensemble. In contrast, the Boltzmann entropy ($S_{B}$) is conventionally defined as $S_{B}=k_{B}\ln\omega$, where
@@
and $\epsilon$ is an absolute energy tolerance, as visualized in Figure 1a.
Physically, the tolerance appears because no system is perfectly isolated, and some amount of perturbation or broadening in the energy occurs due to the system’s interactions with its surroundings (Figure 1b), leading to an “energy window”. See, e.g., Pathria15, who states that “almost every system has some contact with its surroundings, however little it may be; as a result, its energy cannot be defined sharply.” In his work, Gibbs did not comment on the origin of the energy tolerance, simply stating that a Gaussian spread of energy could be considered.1 However, Mayer and Mayer, as well as Landau and Lifshitz provided physical intuition into the energy tolerance using quantum mechanics, whereby Heisenberg’s uncertainty principle requires that $\sigma_{E}\Delta t\geq\frac{\hbar}{2}$, indicating that if $\sigma_{E}$, the uncertainty in energy equals zero, the system would have to remain isolated in the considered energy level for an infinite amount of time, i.e., $\Delta t\rightarrow\infty$.43, 44 In real-world scenarios, however, since the observational time period, i.e., $\Delta t$, is finite, $\sigma_{E}$ also needs to be finite. Note that, in this work, $\Delta E$ represents the difference in energy between two consecutive eigenstates, whereas $\sigma_{E}$ represents the uncertainty in energy of the system, i.e., the energy window.
From Eqs. (3)-(4) and (5)-(6), one can infer that
@@
Even for a discrete spectrum, although direct enumeration is often used as a simplification 45, Eq. (7) is preferred 15, 44, as it allows the correct definition of entropy for nondegenerate quantum systems 46, 47. Indeed, we argue that entropy should account for the uncertainty of a state not only in terms of any degeneracies but also in terms of any perturbations in energy due to the quantum-mechanical restriction on knowing the energy of a state exactly over finite periods of time. Thus, we make a distinction between “degeneracy entropy”, which only accounts for the uncertainty in a microstate due to several microstates possessing exactly the same energy, and the Boltzmann/Gibbs/proposed entropies, which also include the uncertainty due to the allowable spread in the energy, as shown in Figure 2. We show that the degeneracy entropy fails to capture the correct thermodynamic behavior even for most discrete systems, except in the thermodynamic limit in one case.
Typically, $\epsilon$ has been interpreted to be a fixed but small constant in Eq. (7), and this has led to inconsistent results for simple systems with unbounded energy levels, e.g., negative temperatures for an ideal gas with one degree of freedom or infinite temperature for a quantum harmonic oscillator and an ideal gas with two degrees of freedom. We suggest here that instead of using an absolute energy tolerance, the energy tolerance in the microcanonical ensemble should be defined relative to the system’s energy (Figure 1c). Indeed, a system with 1 J of energy and with 1 eV = 1.602 $\times$ 10-19 J of energy should not use the same tolerance in the microcanonical ensemble. This is because the higher the energy of a system, there exists a possibility of greater interaction with its surroundings (as acknowledged by Landau and Lifshitz44) or higher levels of internal perturbations modifying the energy of the system from the calculated energy levels (as remarked by Mayer and Mayer43), causing more broadening or perturbation in the system’s energy.
Now, in principle, a single eigenstate (stationary state) of a system exhibits zero uncertainty in energy, i.e., $\sigma_{E}=0$. However, this is only when it is undisturbed for an infinite amount of time. In practice, the system’s wavefunction would necessarily be entangled with that of its surroundings, involving some level of thermalization. The thermalization time of a quantum system goes as $\frac{h}{E_{\alpha}-E_{\beta}}$ where $\alpha$ and $\beta$ denote different eigenstates, as the eigenstates involved dephase over this timescale.48 Although an upper limit for the thermalization time of many-body, interacting, and perfectly isolated systems is the Heisenberg time, i.e., when $E_{\alpha}-E_{\beta}$ equals the level spacing, $\Delta E$,49 a lower limit for small, noninteracting, and weakly coupled systems could be obtained by considering $E_{\alpha}$ as the eigenstate closest to the mean energy of the system ($E$) and $E_{\beta}$ as the ground state, so that $\Delta t\sim\frac{h}{E}$. For this lower limit of thermalization time, which is inversely proportional to the system’s energy, the uncertainty in energy would be proportional to the system’s energy, as per Heisenberg’s uncertainty principle.
Therefore, we propose that the constraint inside the delta function should be expressed in terms of the relative error of $H_{r}=\frac{H}{E}$ from 1. In terms of the phase space integrals or Hilbert space summations, doing so amounts to applying a constraint on $H$ without using up a degree of freedom. Hence, classically, one should write
@@
where $\epsilon_{r}$ is a relative error tolerance, and is a very small (dimensionless) number independent of $E$. Now, even though $\epsilon_{r}$ affords larger deviations of $H$ from $E$, these deviations are still arbitrarily small due to the former’s small value. Further, as per the above discussion, $\epsilon_{r}E$ represents the uncertainty in knowing the energy of the system exactly, i.e., $\sigma_{E}$, even for a single eigenstate, due to the quantum-mechanical uncertainties introduced by the environment (see Figure 2). Thus, $\epsilon_{r}E=\sigma_{E}$ can be lower than even the energy difference between two successive eigenstates ($\Delta E$), unlike what has been assumed previously ($\sigma_{E}\sim O(\Delta E$)).50
Figure 2: Difference between the proposed, Boltzmann, Gibbs, and degeneracy entropies in terms of which energy levels are included in the calculation of the microcanonical partition function, $w$, denoted using double-ended arrows. $E_{0}$ denotes the ground-state energy, $E_{1}$ the first excited-state energy, and so on. $E_{\max}$ denotes the maximum-allowable energy, which is a finite number for a bounded spectrum, and infinity for an unbounded spectrum. States with energy less than or equal to $E$ are allowed by Gibbs entropy, states with a fixed tolerance $\epsilon$ around $E$, typically of the order of $\Delta E$ are allowed by Boltzmann entropy, states only at energy $E$ are included in $w$ for the degeneracy entropy, and states with a relative-energy tolerance $\epsilon_{r}E$ around $E$ are included in the proposed entropy.
Accordingly, in Eqs. (8) and (9), we use the subscript $r$ to denote the use of the relative deviation from the desired energy. It can be shown that
$\omega_{r}=\epsilon_{r}E\frac{\partial\Omega}{\partial E}$ (10)
as indicated in Figure 1c, in the high-temperature limit, when the states form a continuum. At low temperatures/energies, however, the discreteness of the energy levels implies that
$\omega_{r}=\epsilon_{r}E\frac{\Delta\Omega}{\Delta E}$ (11)
where $\Delta\Omega$ represents the change in the number of states as the energy changes from the present eigenstate to the immediate next one by $\Delta E$. Thus, $\Delta\Omega=1$ for a non-degenerate state, whereas $\Delta\Omega>1$ for degenerate states. In this work, we consider examples of both cases—the single quantum harmonic oscillator and 1D quantum particle in a box are examples where all energy levels are non-degenerate, whereas the two-level system involves degenerate energy levels. In all the examples, we use a forward difference to evaluate $\Delta E$ and $\Delta\Omega$ and found thermodynamically consistent results. We note that apart from issues in defining temperature using the Boltzmann entropy, there are also problems related to the chemical potential, as discussed by Corti et al.6 They remarked that a backward difference is more suitable to evaluate the chemical potential for small systems using the original Boltzmann entropy in the microcanonical ensemble.6 Future work can therefore explore the use of backward differences to evaluate $E$ vs. $T$ relationships, as well as both forward/backward differences for evaluating the chemical potential, using the proposed entropy definition.
We note that there is some precedent in the literature for adopting an energy-dependent tolerance 47, 46. However, in both these studies, the energy tolerance was assumed to be proportional to $\sqrt{C_{C}T_{C}^{2}}$, where $C_{C}=\frac{\partial E}{\partial T_{C}}$ is the canonical heat capacity of the system and $T_{C}$ denotes its canonical temperature. Thus, the energy tolerance depends upon thermodynamic variables ($T_{C}$ and $C_{C}$) not directly defined in the micrononical ensemble, leading to a forced equivalence with the canonical ensemble. In fact, in the Supporting Information, we show that the energy-dependent energy tolerance proposed in refs. 46 and 47 leads to the incorrect equation of state for an isolated quantum harmonic oscillator, unlike the energy tolerance definition proposed in this work. It is also noteworthy that the equations proposed here do not change the definition of the density of states, which is still $\frac{\partial\Omega}{\partial E}$, such that the Laplace transform-based relationship between the canonical partition function and the density of states still holds6. Similarly, the phase space density is unchanged (as it is multiplied everywhere in the phase space by the fixed energy $E$, a constant factor that disappears upon normalization), such that Liouville’s theorem still holds in a manner similar to that for the original Boltzmann entropy.
We now apply the new approach for calculating the microcanonical relative-energy-constrained Boltzmann entropy, $S_{B_{r}}=k_{B}\ln\omega_{r}$, and the corresponding temperature as either
@@
Table 1: Comparison of the predictions of the degeneracy, Gibbs, Boltzmann, and proposed entropies in the microcanonical ensemble with the corresponding predictions in the canonical ensemble.
System $\downarrow$ Entropy $\rightarrow$ Degeneracy Gibbs Boltzmann Proposed Canonical
3D ideal gas (high $T$) not defined $E=\frac{3N}{2}k_{B}T_{G}$ $E=\left(\frac{3N}{2}-1\right)k_{B}T_{B}$ $E=\frac{3N}{2}k_{B}T_{B_{r}}$ $E=\frac{3N}{2}k_{B}T_{C}$
Harmonic oscillator (high $T$) $T_{D}\rightarrow\infty$ $E=k_{B}T_{G}-\frac{h\nu}{2}$ $T_{B}\rightarrow\infty$ $E=k_{B}T_{B_{r}}$ $E=k_{B}T_{C}$
Harmonic oscillator (any $T$) [^1] Note that $E_{0}=\frac{1}{2}h\nu$ and $\theta=\frac{h\nu}{k_{B}}$ for the harmonic oscillator. $T_{D}\rightarrow\infty$ $\frac{E-E_{0}}{h\nu}=\frac{3-\exp\left(\frac{\theta}{T_{G}}\right)}{2\exp\left(\frac{\theta}{T_{G}}\right)-2}$ $T_{B}\rightarrow\infty$ $\frac{E-E_{0}}{h\nu}=\frac{1}{\exp\left(\frac{\theta}{T_{B_{r}}}\right)-1}$ $\frac{E-E_{0}}{h\nu}=\frac{1}{\exp\left(\frac{\theta}{T_{C}}\right)-1}$
Particle in a 1D box (high $T$) $T_{D}\rightarrow\infty$ $E=\frac{k_{B}T_{G}}{2}$ $E=-\frac{k_{B}T_{B}}{2}$ $E=\frac{k_{B}T_{B_{r}}}{2}$ $E=\frac{k_{B}T_{C}}{2}$
[…]
Negative temperatures? Yes No Yes Yes Yes
First, we consider a $d$-dimensional monatomic ideal gas consisting of $N$ atoms of mass $m$ each, occupying a total volume $V$. This system has unbounded energy levels (since there is no upper limit on the classical kinetic energy of a particle) and thus should not have negative absolute temperature, as correctly argued by Dunkel and Hilbert5. The volume of the phase space for this system is 20
$\Omega=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}+1\right)}E^{\frac{dN}{2}}.$ (14)
@@
$\omega=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}\right)}E^{\frac{dN}{2}-1}\epsilon.$ (15)
@@
Note that, for this system, the degeneracy entropy is not defined as the continuum of states implies an uncountable number of precise states. Nevertheless, the Gibbs temperature ($T_{G}$) can be calculated using the equation $\frac{1}{T_{G}}=\frac{\partial S_{G}}{\partial E}$ to obtain $E=\frac{dN}{2}k_{B}T_{G}$. However, using the analogous equation for Boltzmann temperature ($T_{B}$), assuming a fixed energy tolerance $\epsilon$, leads to $E=\left(\frac{dN}{2}-1\right)k_{B}T_{B}$. Note that if $dN=1$, i.e., the system has one degree of freedom, one obtains negative temperatures, whereas if $dN=2$, i.e., the system has two degrees of freedom, one gets infinite temperature 5. (In these cases as well, the center of mass velocity is zero as configurations, wherein the particle moves in both positive and negative directions along a given translational degree of freedom, are included in the phase space average.) This is a serious shortcoming of the Boltzmann entropy, as pointed out by Dunkel and Hilbert 5, as one is led to predict negative/infinite temperature for a system whose energy is unbounded. Considering a fractional energy tolerance, one however obtains $\omega_{r}=\frac{V^{N}}{h^{dN}N!}\frac{(2\pi m)^{\frac{dN}{2}}}{\Gamma\left(\frac{dN}{2}\right)}E^{\frac{dN}{2}}\epsilon_{r}$, which combined with Eq. (12) leads to
$E=\frac{dN}{2}k_{B}T_{B_{r}}.$ (16)
The above equation not only disallows negative/infinite temperatures for an ideal gas but fixes the disagreement between the energy of a three-dimensional ideal gas as traditionally calculated in the canonical $\left(E=\frac{3N}{2}k_{B}T\right)$ and microcanonical $\left(E=\left(\frac{3N}{2}-1\right)k_{B}T\right)$ ensembles (see Table 1)! Although there is no requirement in statistical mechanics for the predictions from various ensembles to match51, similar predictions of the canonical and microcanonical ensemble fall out of the analysis done here using the relative-energy-constrained Boltzmann entropy. This may not always be the case, as shown later in this work for the case of a two-level system. Nevertheless, the relative-energy-constrained Boltzmann entropy eliminates the possibility of negative/infinite temperatures for systems in which there is no maximum limit on the allowed energies. Moreover, the proposed entropy fixes the disagreement between the ideal gas temperature based on the inverse kinetic energy as obtained from the microcanonical ensemble and that from the zeroth law of thermodynamics.6